# Kinetic energy in random recurrent neural networks

### v1

@@ Abstract
The relationship between unstable fixed points and chaotic dynamics in high-dimensional neural dynamics remains elusive. In this work, we investigate the kinetic energy distribution of random recurrent neural networks by combining dynamical mean-field theory with extensive numerical simulations. We find that the average kinetic energy shifts continuously from zero to a positive value at a critical value of coupling variance (synaptic gain), with a power-law behavior close to the critical point. The steady-state activity distribution is further calculated by the theory and compared with simulations on finite-size systems. This study provides a first step toward understanding the landscape of kinetic energy, which may reflect the structure of phase space for the non-equilibrium dynamics.
@@ I Introduction
Recurrent neural networks (RNNs) with random couplings exhibit a continuous dynamical phase transition from a trivial single fixed point to non-equilibrium chaotic high-dimensional dynamics [1, 2, 3]. The dynamical phase transition shows self-organized criticality via tuning only the variance of synaptic gain despite different forms of synaptic connectivity, which offers computational advantage for reservoir computing [4, 5, 6] and toy models for better understanding population computation in real brain circuits [7, 8, 9, 10]. However, understanding the nature of RNNs remains incomplete. A hot current debate is about whether the unstable fixed points after the chaos transition are related to the chaotic dynamics. Recent works argued that the fixed point ensemble provides transiently attractive landmarks that guide the chaotic flow [11, 12]. Moreover, fixed points and dynamics are confined to separate shells in the phase space [11, 13]. Nevertheless, other works challenged this viewpoint and argued that the high-dimensional chaos can not be inferred from these equilibria, showing theoretical evidence in a mixture model of conservative and non-conservative forces [14].
Inspired by searching for fixed points (equilibria) in RNNs [15], we formulated the fixed point distribution as an optimization problem of minimizing kinetic energy of the dynamics [3]. To explore the relationship between the equilibria and chaotic dynamics, one first should study the landscape of kinetic energy itself. This is the main focus of the paper. By applying dynamical mean-field theory (DMFT) [16, 14], we identify a continuous transition for the stationary (in the long time limit) kinetic energy and determine the scaling exponent for the critical behavior close to the transition point. The theory is corroborated by numerical simulations. Moreover, the single time probability distribution of the chaotic activity is also derived and calculated by the DMFT, supported by simulations in a finite-size system. As kinetic energy grows, this distribution gets broadened as well. We expect this theoretical result to finally lead to new insights toward the debate.
@@ II Model setting and DMFT
We consider a recurrent neural network composed of $N$ interacting units whose dynamics evolve according to the following equation:
@@
where $x_{i}(t)$ denotes the synaptic current (transformed to activity through $\phi(\cdot)$) of neuron $i$ at time $t$, $\phi(x)$ is a nonlinear activation function [in this work we use $\text{tanh}(\cdot)$], and $J_{ij}$ are synaptic weights drawn independently from a Gaussian distribution with zero mean and variance $\frac{g^{2}}{N}$. Here, $g$ is also referred to as the synaptic gain, which controls the overall level of recurrent synaptic strength in the network. When $g$ crosses a threshold of $g_{c}=1$, a dynamical phase transition is triggered from a fixed point of null activity to chaos [1]. The entire system is spontaneously driven by a fragile state slightly above $g_{c}$ (Fig. 1).
Figure 1: Sketch of network connectivity and dynamics. (a) Schematic illustration of the RNN connectivity structure. (b) Sample trajectories of five neurons for coupling strengths $g=0.5,1.05,2$.
In the limit of a large network size ($N\rightarrow\infty$), we apply the framework of dynamical mean-field theory (DMFT) to reduce the high-dimensional system to an effective single-unit description. Under DMFT, the dynamics of a representative neuron are governed by the following stochastic differential equation:
@@
where $\eta(t)$ is a Gaussian random process of zero mean that encapsulates the cumulative effect of the recurrent inputs from the rest of the network. This mean-field process can also be derived by applying the cavity method in a self-consistent way [2]. In the original network, the input current to neuron $i$ is given by $\sum_{j}J_{ij}\phi(x_{j}(t))$. Since the weights $J_{ij}$ are drawn independently from a Gaussian distribution, independent of neural activity as well, the total input can be approximated as a Gaussian random variable due to the central limit theorem. Crucially, the effective noise $\eta(t)$ is not white, but temporally correlated. The second-order statistics are thus determined self-consistently by the network’s own dynamics:
@@
which means that the noise $\eta(t)$, although modeled as an external Gaussian field in the reduced mean-field description, actually emerges from the collective activity of the network itself.
@@
This leads to the explicit form by applying Price’s theorem [Price-1958]:
@@ III Kinetic energy analysis
In this section, we present the results of our numerical simulations and theoretical predictions for the average kinetic energy in random RNNs. We also derive the scaling law of kinetic energy in the critical regime where $g\rightarrow 1^{+}$. Furthermore, we examine the stationary distribution of RNN activity obtained from DMFT and analyze the arc length of neural trajectories, which provides insights into the temporal evolution of the network dynamics.
@@ III.1 Average kinetic energy
Starting from Eq. (2), multiplying both sides and taking the long-time limit, we obtain:
@@
where the Fourier transformation is applied (Technical details can be found in Appendix A). According to Eq. (7), which relates $\partial_{\tau}^{2}\Delta_{\tau}$ to $C_{\tau}$, the average kinetic energy ($\frac{1}{2}\mathbf{v}^{2}$, where $\mathbf{v}$ is the velocity) can be written as $\Gamma(\Delta;\Delta_{0})$ in the TTI limit of $\Gamma_{tt^{\prime}}$:
@@
$\lim\limits_{\tau\rightarrow\infty}\varGamma(\tau)=0.$ (14)
@@
which allows us to compute the DMFT prediction for the average kinetic energy. We use a Python-based numerical solver to find the solution of $\mathcal{F}(\Delta_{0};\Delta_{0})=0$, and then substitute the resulting $\Delta_{0}$ back into Eq. (13). During the numerical procedure, we set a convergence threshold of $10^{-10}$, and ensure that each solution satisfies $\mathcal{F}(\Delta_{0};\Delta_{0})<10^{-5}$. We repeat the root-finding process 100 times and compute the average value of $\Gamma(\Delta_{0})$ (denoted as $\Gamma_{0}$ below), with the standard deviation shown as an error bar in the plot.
Figure 2 compares the average kinetic energy obtained from direct simulation of Eq. (1) with the DMFT prediction from Eq. (13). The left panel shows results for a wide range of coupling strengths $0.8\leq g\leq 1.5$, while the right panel zooms in the near-critical region $0.9\leq g\leq 1.2$. We observe excellent agreement between the simulation results and the DMFT theory. Moreover, the fluctuations in the simulated data decrease as the network size increases, as expected.
Figure 2: The average kinetic energy of the system versus synaptic gain $g$. (a) Results for $0.8\leq g\leq 1.5$; (b) An enlarged view of the critical regime $0.9\leq g\leq 1.2$. Simulations were performed using networks of size $N=1000$, $5000$, and $10000$, with error bars computed from 100 independent realizations. The DMFT result is obtained by numerically solving $\mathcal{F}(\Delta_{0};\Delta_{0})=0$ for $\Delta_{0}$, and substituting the solution into Eq. (13). Error bars for the DMFT prediction are computed from 100 numerical evaluations.
@@ III.2 Critical behavior of kinetic energy at $g\rightarrow 1^{+}$
Figure 3: Comparison between the theoretical prediction [Eq. (18) or Eq. (20)] in the limit $g\to 1^{+}$ and the numerical results from DMFT. (a) The dependence of $\Delta_{0}$ on $\sigma$; (b) $\Gamma_{0}$ obtained by substituting $\Delta_{0}$ into the analytical expression [Eq. (19)]. The DMFT error bar is estimated from 100 independent numerical evaluations. (c) Enlarged view of the comparison close to the critical point in log-log scale. Note that the mean value of $\Delta_{0}$ is inserted into Eq. (19).
@@
Using Eq. (18), we derive immediately the asymptotic behavior of the kinetic energy near the critical point $g\rightarrow 1^{+}$:
@@
The detailed derivation of these results is provided in Appendix B. As shown in Fig. 3, the theoretical prediction of the kinetic energy in the critical regime is in excellent agreement with the numerical solution of the DMFT equations. This confirms that the scaling law of the kinetic energy near the criticality follows roughly a cubic behavior, with the theoretical scaling exponent equal to $3$.
III.3 Single-time probability distribution of stationary activity
@@
$x(t)=\int_{-\infty}^{t}\mathrm{d}se^{-(t-s)}\eta(s).$ (22)
The variance is determined self-consistently by solving the single-unit DMFT equations [Eq. 2, see also Appendix C].
Figure 4: Distribution of $x(t)$ at steady states for different coupling strengths of $g=1.1,1.3,1.5$. Histograms represent numerical simulation results of the RNN dynamics ($N=5000$), while dashed lines correspond to the Gaussian distributions obtained from DMFT calculations.
@@
In Fig. 5, the dynamical trajectories of the RNN and the corresponding trajectories from the DMFT equations are shown for different values of $g$. As $g$ increases, the fluctuations in the trajectories become more pronounced. The algorithmic steps for solving the DMFT equation [Eq. (2)] can be found in Appendix. C.

### v2

@@ Abstract
High-dimensional chaotic dynamics can emerge in a large random recurrent neural network when the synaptic gain crosses a threshold. Recent works showed that the kinetic energy of neural activity links the chaotic dynamics and the supporting unstable fixed points (equilibria) in the phase space. Here, we investigate the kinetic-energy-centric properties of random recurrent neural networks by combining dynamical mean-field theory with extensive numerical simulations. We find that the average kinetic energy shifts continuously from zero to a positive value at a critical value of coupling variance (synaptic gain) and exhibits a cubic scaling behavior near the critical point from above. This scaling behavior is supported by numerical simulations and provides a quantitative characterization of how fast the dynamics change during the onset of chaos. The steady-state activity distribution is further calculated by the theory and compared with simulations on finite-size systems from the kinetic-energy optimization perspective as well. The activity distribution is also analyzed in a geometric angle, establishing a relationship between the original chaotic dynamics and the gradient dynamics of the kinetic energy. The trajectory length on the chaotic manifold can be derived from the stationary kinetic energy, and the associated stationary behavior is analyzed as well. This study provides a kinetic-energy-centric route toward understanding the dynamics landscape of recurrent neural networks, which may provide insights for reservoir computing and even for internal synaptic learning.
@@ I Introduction
High-dimensional recurrent neural networks (RNNs) with random synaptic couplings provide a minimal yet powerful theoretical framework for understanding nonequilibrium dynamics, emergent computation, and collective fluctuations in complex systems. Increasing synaptic gain induces a continuous transition from a stable fixed point to high-dimensional chaos [1]. This intriguing collective behavior makes the random RNN a paradigmatic model in statistical physics, neuroscience, and machine learning. A standard tool for analyzing this chaos transition is dynamical mean-field theory (DMFT) [2, 3, 4, 5, 6, 7], which offers an exact description of the non-equilibrium steady dynamics in the thermodynamic limit.
@@
RNNs with random couplings exhibit a continuous dynamics phase transition slightly above which the dynamics are extremely slow but sensitive to external perbubations [1, 5, 8]. The dynamics phase transition thus shows self-organized criticality via tuning only the variance of synaptic coupling despite different forms of synaptic connectivity, which offers computational advantage for reservoir computing [9, 10, 11, 12]. The model also provides a theoretical basis for better understanding population computation in real brain circuits [13, 14, 15, 16, 17, 18]. However, understanding the nature of chaotic dynamics in RNNs remains incomplete. A hot current debate is about whether the unstable fixed points after the chaos transition are related to the chaotic dynamics. Recent works argued that the fixed point ensemble provides transiently attractive landmarks that guide the chaotic flow [19, 20]. Moreover, fixed points and dynamics are confined to separate shells in the phase space [19, 21, 22]. Nevertheless, other works challenged this viewpoint and argued that the high-dimensional chaos can not be inferred from these equilibria, showing theoretical evidence in a mixture model of conservative and non-conservative forces [23].
Inspired by searching for fixed points (equilibria) in RNNs [24], we formulated the fixed point distribution as an optimization problem of minimizing kinetic energy of the dynamics in a recent work [8]. How the kinetic energy varies with the synaptic gain and the associated impact on the dynamics on the chaotic attractor remain unclear. We will address this question in the current work. By applying dynamical mean-field theory [2, 23], we find that at the chaos transition point, the stationary (in the long time limit) kinetic energy becomes positive in a continuous way, and we determine the scaling exponent for the critical behavior close to the transition point. A cubic growth is identified. Interestingly, this behavior is shared with the linear dimensionality growth [25] as well as the topological complexity growth [19], differently from the quadratic growth of the maximum Lyapunov exponent during the onset of chaos [1, 26]. The theory is corroborated by numerical simulations. Moreover, the single-time probability distribution of the chaotic activity is also derived and calculated by the DMFT, supported by simulations in a finite-size system and compared with the gradient dynamics under the kinetic energy as well. The chaotic steady state is mapped to stationary dynamics of a Langevin-type gradient system with an effective temperature fixed by kinetic-energy matching. Both systems share the same distribution of synaptic currents, yet their activities are located at different shells with a spherical rotation (revealed in finite system simulations). Whether this holds in infinite-sized systems remains to be explored in the future. Finally, we show that the arc length of chaotic trajectories grows linearly with a rate determined by the stationary kinetic energy. We expect this work will improve our theoretical understanding of neural computation at the onset of chaos and far beyond the edge of chaos.
@@ II Model setting and DMFT
This section reviews the standard dynamical mean-field theory for random recurrent neural networks, following Refs.[1, 2, 27, 28], since the kinetic energy relies on the DMFT framework. The part particularly related to the derivation of kinetic energy and its asymptotic behavior is our contribution in the current work and will be mentioned in the next section. We consider a recurrent neural network composed of $N$ interacting units whose dynamics evolve according to the following equation:
@@
where $x_{i}(t)$ denotes the synaptic current [transformed to activity through $\phi(\cdot)$] of neuron $i$ at time $t$, $\phi(x)$ is a nonlinear activation function [in this work we use $\text{tanh}(\cdot)$], and $J_{ij}$ are synaptic weights drawn independently from a Gaussian distribution with zero mean and variance $\frac{g^{2}}{N}$. Here, $g$ is also referred to as the synaptic gain, which controls the overall level of recurrent synaptic strength in the network. When $g$ crosses a threshold of $g_{c}=1$, a dynamics phase transition is triggered from a fixed point of null activity to chaos [1]. Slightly above the critical coupling strength $g_{c}$, the system enters an unstable regime, where the dynamics become highly sensitive to small perturbations and spontaneously evolve into a dynamics regime of chaotic fluctuations (Fig. 1).
Figure 1: Sketch of network connectivity and dynamics. (a) Schematic illustration of the RNN connectivity structure. (b) Sample trajectories of five neurons for coupling strengths $g=0.5,1.05,2$ in a neural pool of N=5  [^0] N=5\,000.
In the limit of a large network size ($N\rightarrow\infty$), we apply the framework of dynamical mean-field theory to reduce the high-dimensional system to an effective single-unit description. Under DMFT, the dynamics of a representative neuron are governed by the following stochastic differential equation:
@@
where $\eta(t)$ is a Gaussian random process of zero mean that encapsulates the cumulative effect of the recurrent inputs from the rest of the network. Crucially, the effective noise $\eta(t)$ is not white, but temporally correlated. This mean-field process can be derived by applying the cavity method in a self-consistent way [5]. In the original network, the input current to neuron $i$ is given by $\sum_{j}J_{ij}\phi(x_{j}(t))$. Within the DMFT framework, the input is modeled as a Gaussian process whose statistics are self-consistently determined by one-dimensional effective dynamics. In essence, the Gaussian description emerges from a cavity-type argument in the large-network limit or from a path-integral formalism (see technical details in a lecture note [5]). The second-order statistics are thus determined self-consistently by the network’s own dynamics:
@@
where $\phi(t)\equiv\phi(x(t))$. Note that the noise $\eta(t)$, although modeled as an external Gaussian field in the reduced mean-field description, actually emerges from the collective activity of the network itself.
@@
This leads to the explicit form by applying Price’s theorem [29]:
@@ III Kinetic energy analysis
In this section, we present the results of our numerical simulations and theoretical predictions for the average kinetic energy in random RNNs. We also derive the scaling law of kinetic energy in the critical regime where $g\rightarrow 1^{+}$. Furthermore, we examine the stationary distribution of RNN activity obtained from DMFT in comparison with that obtained through gradient dynamics on kinetic energy keeping an equivalent temperature, and their geometric locations are also compared, and finally we analyze the arc length of neural trajectories, which depends on the steady state kinetic energy and provides insights into the temporal evolution of the network dynamics in the steady state regime.
@@ III.1 Average kinetic energy
Starting from Eq. (2), multiplying both sides and taking the long-time limit, we obtain the speed correlation:
@@
where $\tau=t-t^{\prime}$, and the Fourier transformation is applied (Technical details can be found in Appendix A). This speed correlation was first explored in a different context [23]. Here we give an alternative derivation of $\Gamma_{tt^{\prime}}=-\partial_{\tau}^{2}\Delta_{\tau}$ by Fourier transformation. This correlation can be linked to the kinetic energy studied earlier in Ref. [8]. The kinetic energy is defined as the $\ell_{2}$ norm of the state velocity, namely $\frac{1}{2}\mathbf{v}^{2}$, where $\mathbf{v}$ is the velocity. It quantifies how fast the system moves through its high-dimensional state space and therefore provides a trajectory-level characterization of the dynamics. Therefore, the kinetic energy serves as a physical reflection of the dynamics landscape, which we shall explore in detail later. According to Eq. (7), which relates $\partial_{\tau}^{2}\Delta_{\tau}$ to $C_{\tau}$, the average kinetic energy can be written as $\Gamma(\Delta;\Delta_{0})$ in the TTI limit of $\Gamma_{tt^{\prime}}$:
@@
$\lim\limits_{\tau\rightarrow\infty}\Gamma(\tau)=0.$ (14)
@@
which allows us to compute the DMFT prediction for the average kinetic energy. We use a Python-based numerical solver to find the solution of $\mathcal{F}(\Delta_{0};\Delta_{0})=0$, and then substitute the resulting $\Delta_{0}$ back into Eq. (13). During the numerical procedure, we set a convergence threshold of $10^{-10}$, and ensure that each solution satisfies $\mathcal{F}(\Delta_{0};\Delta_{0})<10^{-5}$. We repeat the root-finding process $100$ times and compute the average value of $\Gamma(\Delta_{0})$ (denoted as $\Gamma_{0}$ below), with the standard deviation shown as shaded regions in the plot.
Figure 2 compares the average kinetic energy obtained from direct simulation of Eq. (1) with the DMFT prediction from Eq. (13). The left panel shows results for a wide range of coupling strengths $0.8\leq g\leq 1.5$, while the right panel zooms in on the near-critical region $0.9\leq g\leq 1.2$. We observe excellent agreement between the simulation results and the DMFT theory. Moreover, the fluctuations in the simulated data decrease as the network size increases, as expected. Here the shaded error of the DMFT theory in Fig 2 (b) stems mainly from the integral in Eq. (13). We employed $10^{7}$ Monte Carlo samples to compute this integral, resulting in an error of the order $10^{-3}$.
Figure 2: The average kinetic energy of the system versus synaptic gain $g$. (a) Results for $0.8\leq g\leq 1.5$; (b) An enlarged view of the critical regime $0.9\leq g\leq 1.2$. Simulations were performed using networks of size $N=1000$, $5000$, and $10000$, with standard deviations computed from $100$ independent realizations. The DMFT result is obtained by numerically solving $\mathcal{F}(\Delta_{0};\Delta_{0})=0$ for $\Delta_{0}$, and substituting the solution into Eq. (13). Standard deviations for the DMFT prediction are computed from $100$ numerical evaluations (see a detailed explanation in the main text).
@@ III.2 Critical behavior of kinetic energy at $g\rightarrow 1^{+}$
Here, we obtain $\Delta_{0}$ following the method presented in Ref. [2], with a slightly different form of dynamics. We give more details in Appendix B, which is absent in Ref. [2]. Equation (18) will be used below to determine the scaling behavior of the kinetic energy.
Figure 3: Comparison between the theoretical prediction [Eq. (18) or Eq. (20)] in the limit $g\to 1^{+}$ and the numerical results from DMFT [solutions of $\mathcal{F}(\Delta_{0};\Delta_{0})=0$ or Eq. (13)]. (a) The dependence of $\Delta_{0}$ on $\sigma$; (b) $\Gamma_{0}$ obtained by substituting $\Delta_{0}$ into the analytical expression [Eq. (19)]. The DMFT shaded error is estimated from $200$ independent numerical evaluations via Eq. (13). (c) Enlarged view of the comparison close to the critical point in log-log scale. Note that the mean value of $\Delta_{0}$ is inserted into Eq. (19).
@@
Using Eq. (18), we derive the asymptotic behavior of the kinetic energy near the critical point $g\rightarrow 1^{+}$:
@@
The detailed derivation of these results is provided in Appendix B. From the above derivations in this section, we obtain an analytical result for the kinetic energy, which scales with $\sigma$ as a cubic growth law when $g\to g_{c}$ from above. As shown in Fig. 3, the theoretical prediction of the kinetic energy in the critical regime is in excellent agreement with the numerical solution of the DMFT equations. The numerical solution for $\Delta_{0}$ obtained via numerical root finding exhibits negligible errors [Fig. 3 (a)]. In contrast, the DMFT results, due to the integral in Eq. (13), display an error of the order $10^{-4}$ in Fig. 3 (b). Our theoretical result [Eq. (20)] begins to deviate from the DMFT prediction as the gain factor $g$ increases, because the expansion is only valid in the vicinity of the critical point [Fig. 3 (c)]. This confirms that the scaling law of the kinetic energy near the criticality follows roughly a cubic behavior, with the theoretical scaling exponent equal to $3$.
The nature of neural computation at the onset of chaos can also be understood by other physical quantities. The maximal Lyapunov exponent quantifies the exponential separation rate of nearby trajectories and has been shown to grow quadratically with $\sigma$ [1, 26]. By contrast, the linear dimensionality of chaotic attractors bears a cubic growth slightly above the transition [25], and the topological complexity of unstable fixed points has also a cubic growth [19]. Interestingly, the kinetic energy also displays a cubic growth just above the critical point. They all may have a common root in the dynamics and fixed-point landscape. It is promising in future works to elucidate this hidden relationship.
III.3 Single-time probability distribution and geometric structure of stationary activity
@@
$x(t)=\int_{0}^{t}\mathrm{d}se^{-(t-s)}\eta(s).$ (22)
The variance is determined self-consistently by solving the single-unit DMFT equations [Eq. (2), see also Appendix C].
Figure 4: Probability distribution of stationary activity. (a) Distribution of $x(t)$ at steady states for different coupling strengths $g=1.1,1.3,1.5$. Histograms represent numerical simulation results from RNN simulations ($N=5000$), while dashed lines correspond to the Gaussian distributions obtained by DMFT. (b) Mean kinetic energy as a function of temperature ($g=1.5,N=5000$). The blue solid curve represents the steady-state average kinetic energy of the gradient descent (GD) dynamics in Eq. (23). The purple dashed curve shows the corresponding steady-state average for the RNN dynamics in Eq. (1) with the same interaction matrix. The vertical dashed line indicates the effective temperature (mean matched with the RNN counterpart). The mean kinetic energy and the associated error bars (only shown for the matched point) are computed from $200$ statistically independent samples taken in the stationary regime. (c) Comparison between sampled steady state activities of the two dynamics (the GD one at the effective temperature) ($g=1.5,N=5000$).
@@
In Fig. 5, the dynamical trajectories of the RNN and the corresponding trajectories from the DMFT equations are shown for different values of $g$. As $g$ increases, the fluctuations in the trajectories become more pronounced. The algorithmic steps for solving the DMFT equation [Eq. (2)] can be found in Appendix C. This result provides an intuitive understanding about why the distribution of stationary activity gets broadened in Fig. 4 and about why the arc length growth rate increases with $g$ in the next section.
@@
Next, we explore how the stationary activity is geometrically distributed. We consider a kinetic energy optimization, which will be compared with the original non-gradient RNN dynamics. Introduced in an earlier work [8], the kinetic-energy optimization dynamics read
$\dot{\mathbf{x}}=-\nabla_{\mathbf{x}}E(\mathbf{x})+\sqrt{2T}\,\bm{\epsilon}(t),$ (23)
with Gaussian white noise $\bm{\epsilon}(t)$ and kinetic energy
$E(\mathbf{x})=\frac{1}{2}\sum_{i=1}^{N}\left(-x_{i}+\sum_{j=1}^{N}J_{ij}\phi(x_{j})\right)^{2}.$ (24)
By construction, this potential system admits an equilibrium Boltzmann distribution [30]
$P(\mathbf{x})=\frac{1}{Z}\exp\left(-\beta E(\mathbf{x})\right),$ (25)
where $\beta=\frac{1}{T}$ is related to the noise stength in Eq. (23).
We compute the average energy $\langle E\rangle$ as a function of temperature $T$ for the potential system and compare the result with the mean kinetic energy of the original RNN in the stationary state ($E=\frac{1}{2}\Gamma_{0}$), as shown in Fig. 4 (b). This comparison defines an effective temperature $T_{\mathrm{eff}}$ via a bisection procedure, such that the mean kinetic energy of the two dynamics coincides. Importantly, $T_{\mathrm{eff}}$ is not introduced as a free parameter, but determined by the intrinsic fluctuation strength generated by chaotic dynamics.
When the gradient dynamics is simulated at $T_{\mathrm{eff}}$, the resulting stationary single-neuron activity distribution matches that of the original RNN, as demonstrated in Fig. 4 (c). This result shows that, at the level of single-time statistics, the nonequilibrium stationary state of the chaotic RNN can be mapped onto an equilibrium-like ensemble characterized by an effective temperature. From this perspective, the chaotic fluctuations play a role analogous to thermal noise in a corresponding potential system.
Figure 6: Geometric structure of stationary activity for RNN and gradient dynamics at $g=1.5$. (a) Two-dimensional geometric representation of stationary activity samples for system size $N=5000$. Each point indicates a single high-dimensional state projected onto a polar representation with radius $\left\|\phi\right\|$ and angle $\psi$ with the ensemble average center $c_{\rm G}$ or $c_{\rm R}$. Purple (RNN) and blue (GD) clouds correspond to the original RNN dynamics and the associated gradient dynamics, respectively. Solid arrows indicate the mean vectors $\mathbf{c}_{R}$ and $\mathbf{c}_{G}$, while dashed lines mark the realm of the distributions. Inset: Angle $\theta$ between the center vectors of the RNN and gradient dynamics as a function of the number of stationary samples $M$. The samples are obtained by uniformly sampling different time points along a single stationary trajectory for each dynamics, using the same realization of the coupling matrix and identical initial conditions. The red point indicates that M=3  [^0] M=3\,000 is used for the current cloud representation. (b) Probability desity $P(\psi)$ of the angular coordinate $\psi$ for system sizes $N=2000$ and $N=5000$. For each system size, the gradient dynamics are simulated at the corresponding effective temperature, i.e., $T_{\mathrm{eff}}=0.1140~(N=2000)$ and $T_{\mathrm{eff}}=0.1093~(N=5000)$. The two dynamics exhibit distinct angular dispersion and a finite misalignment angle $\theta$ between their centroid directions. (c) Probability density of the radial norm $\left\|\phi\right\|$. All samples are taken from a single long trajectory in the stationary regime in order to probe the structure of the concrete attractor.
An immediate remaining question is whether both dynamics occupy the same part of the phase space. To address this question, we perform the following analysis. We consider the original RNN dynamics (denoted by R) and the associated gradient dynamics (denoted by G), using the same realization of the coupling matrix $\mathbf{J}$ and identical initial conditions. As assumed, the dynamics will cover the entire manifold given an infinite simulation time. For the gradient dynamics, the temperature is chosen to be equal to the effective temperature extracted from the RNN dynamics [see Fig. 4 (b)].
Starting from the common initial state, each system is evolved in time until entering the stationary regime, as verified by the changes of macroscopic observables such as kinetic energy. In the stationary regime, we record system states every fixed time interval, thereby reducing temporal correlations between successive samples. In total, we collect M=3  [^0] M=3\,000 stationary samples from each dynamics, yielding two ensembles $\{\phi(\mathbf{x})_{R}^{\nu}\}$ and $\{\phi(\mathbf{x})_{G}^{\nu}\}$, $\nu=1,2,\cdots,M$. From each ensemble, we compute the centroid vector:
$\mathbf{c}_{R}=\big\langle\phi(\mathbf{x})_{R}\big\rangle_{M},\quad\mathbf{c}_{G}=\big\langle\phi(\mathbf{x})_{G}\big\rangle_{M},$ (26)
where $\left\langle\cdot\right\rangle_{M}$ denotes an average over the $M$ stationary samples. For each sample, we then define the polar angle with respect to its corresponding centroid vector as follows [22]
$\psi_{R}^{\nu}=\arccos\left(\frac{\phi(\mathbf{x})_{R}^{\nu}\cdot\mathbf{c}_{R}}{\|\phi(\mathbf{x})_{R}^{\nu}\|\|\mathbf{c}_{R}\|}\right),\quad\psi_{G}^{\nu}=\arccos\left(\frac{\phi(\mathbf{x})_{G}^{\nu}\cdot\mathbf{c}_{G}}{\|\phi(\mathbf{x})_{G}^{\nu}\|\|\mathbf{c}_{G}\|}\right);$ (27)
and also calculate the angle between the two centroid vectors:
$\theta=\arccos\left(\frac{\mathbf{c}_{R}\cdot\mathbf{c}_{G}}{\|\mathbf{c}_{R}\|\|\mathbf{c}_{G}\|}\right).$ (28)
As illustrated in Fig. 6, the stationary states of the two dynamics form dispersed clouds in the phase space, organized around their respective centroids. The angular dispersion observed in Fig. 6 (b) corresponds to the $\Delta{\psi}$ illustrated in Fig. 6 (a), while the activity $\ell_{2}$ norm distribution $\|\phi\|$ is relatively narrow and shown in Fig. 6 (c) corresponding to the difference between the radii of the outer and inner circular envelopes in Fig. 6 (a). Figure 6 (b) shows that the angular distributions differ markedly between the two systems, with a rotation of polar angle $\theta$ in Fig. 6 (a)]. The inset of Fig. 6 (a) shows that even when $M$ is increased up to $2\times 10^{4}$, the angle $\theta$ does not converge to a fixed value but instead exhibits persistent fluctuations over a broad range. Although both dynamics are arranged in a shell-like structure, they are well separated in the polar direction. Whether this geometric property holds in the thermodynamic limit deserves further study in future theoretical works. In the zero temperature limit, the potential dynamics will approach the intrinsic unstable fixed-points of the original RNN [20], but these fixed points repel the dynamics [19]. A recent work studied an inhibition-dominated neural system and found that the dynamical cap always lies inside the region containing the fixed points [22].